Show that the phase difference between displacement and velocity is 90 degree and between the displacement and acceleration is 180 degree

Grade:12th pass

1 Answers

harini
7 years ago
      from the graphs of accelration,velocity and displacemnt(please look for them in your 11th NCERT ;) i couldnt copy, paste the image :/ ) it is evident that velocity lags back displacement by 90 and acceleration leads displacement by 180. you can also verify it using the basic SHM equation x(t)= Acos(wt+@)   for velocity it will be V(t)= -Awsin(wt+@) {by differentitaing displacement equation, this equation of velocity is obtained) similarly differentitiate acceleration , we’ll be getting a(t)= -Aw^2cos(wt+@).this last equation can be again simplified as a(t)= -w^2x(t) clearly depicts acceleration is proportional to displacement but out of phase(180) with displacement and this is the one of the most essential conditions of a particle to be in SHM. 
    Hope it helps :)